# Extension of the CC(;) Formalism to the Electron Attachment and Ionization Potential Equation-of-Motion Coupled-Cluster Frameworks

### v1

@@ Abstract
keywords:
Coupled-Cluster Theory , Electron Attachment and Ionization Potential Equation-of-Motion Coupled-Cluster Methods , Completely Renormalized Coupled-Cluster Approaches , CC(;) Formalism1 Introduction
@@ 2 Theory
This study focuses on the CC(;) approaches aimed at accurately approximating the EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) energetics at small fractions of the computational costs by correcting the results of the EA/IP-EOMCC calculations with all 1/1, all 2-1/2-1, and, optionally, some 3-2/3-2 excitations on top of CCSD for the many-electron correlation effects associated with all or the remaining 3-2/3-2 contributions using the moment expansions defined by Eq.Β (9). In the case of the noniterative corrections to EA-EOMCCSD(2-1) due to 3-2 correlations, which define the EA-CR-EOMCC(2,3) method, the space is spanned by all 1 () and 2-1 () determinants and the corresponding space consists of all 3-2 determinants (). Similarly, the IP-CR-EOMCC(2,3) approach, which corrects the IP-EOMCCSD(2-1) energetics for 3-2 effects, uses a space spanned by all 1 () and 2-1 () determinants and a space consisting of all 3-2 determinants (). In the case of the active-orbital-based EA-CC(t;3) methodology, the space is defined as all and determinants and the subset of 3-2 determinants selected according to the formula , where the index runs over active unoccupied spinorbitals defined by the user, and the space consists of the remaining 3-2 determinants that are not included in the space. Finally, the IP-CC(t;3) approach adopts a space spanned by all , all , and the subset of 3-2 determinants defined as , where the index denotes the user-defined active occupied spinorbitals, and the space is spanned by the remaining 3-2 determinants.
@@
In analogy to the CC(;) methods of Refs.Β [27, 28, 33, 34, 35, 36] and their CR-CC(2,3) and CR-EOMCC(2,3) predecessors, the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches offer considerable savings in the computational effort compared to their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents. This is because the excitation spaces used in the iterative EA/IP-EOMCC steps contain none or only some 3-2/3-2 determinants and the noniterative corrections require fewer CPU operations than a single iteration of EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2). In the next section, we assess how effective the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) methods are in improving the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt computations and recovering the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) energetics. In doing so, we examine two variants of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, termed EA/IP-CR-EOMCC(2,3)X and EA/IP-CC(t;3)X, respectively, where or . Following the naming convention adopted in Refs.Β [31, 32, 33], the methods labeled with subscript use the EN-like energy denominators in Eq.Β (9) when evaluating corrections . The methods using subscript replace the EN form of by its MP counterpart.
@@ 3.1 Computational Details
To assess the performance of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, we applied them, along with the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt methods and their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents, to the ground and valence excited states of the CH, CNC, C2N, SH, N3, and NCO radical species. The results of our EA-EOMCC calculations for CH, CNC, and C2N are reported in Tables 1 and 2, whereas Tables 3 and 4 summarize the IP-EOMCC computations for SH, N3, and NCO. With the exception of SH, all of our EA/IP-EOMCC calculations correspond to adiabatic transitions; calculations for SH were performed at the ground-state equilibrium geometry. The nuclear geometries characterizing the ground and excited states of CH and the ground-state SβH separation in SH were the same as those used in Ref.Β [15], whereas the CNC, C2N, N3, and NCO molecules were described using the linear geometries optimized with the EA (CNC and C2N) or IP (N3 and NCO) variants of the symmetry-adapted-cluster (SAC) configuration interaction (CI) approach [38] abbreviated as SDT- [39] in Ref.Β [18]. All EA-EOMCC calculations for CH, CNC, and C2N reported in this work employed the restricted HartreeβFock (RHF) orbitals of the underlying (CH)+, (CNC)+, and (C2N)+ cationic references, whereas the IP-EOMCC calculations for SH, N3, and NCO used the RHF orbitals of the closed-shell (SH)-, (N3)-, (NCO)- anions. Our calculations for CH employed the aug-cc-pVQZ basis set [40, 41], while for SH, we used the aug-cc-pV(T+d)Z [42] basis for sulfur and aug-cc-pVTZ [40, 41] for hydrogen. Following Ref.Β [18], the CNC, C2N, N3, and NCO radicals were described using the DZP[4s2p1d] basis set [43, 44]. The active orbital spaces used to define the and components of entering the EA/IP-EOMCCSDt and subsequent EA/IP-CC(t;3) calculations were the same as those used in Refs.Β [15, 18]. Thus, the EA-EOMCCSDt and EA-CC(t;3) calculations for CH used an active space consisting of the lowest-energy orbitals unoccupied in (CH)+, whereas the active spaces employed in the analogous calculations for CNC and C2N were spanned by the four orbitals corresponding to the two lowest shells unoccupied in (CNC)+ and (C2N)+. In the case of the IP-EOMCCSDt and IP-CC(t;3) computations for SH, N3 and NCO, the active spaces consisted of the highest-energy shell occupied in the respective parent (SH)-, (N3)-, and (NCO)- anions. The lowest-energy orbitals corresponding to the 1s shells of carbon, oxygen, nitrogen, and sulfur atoms were frozen in all post-RHF steps.
@@ 3.2 Performance of the CC(;)-based EA/IP-EOMCC Methodologies
We begin our discussion by examining the EA-EOMCC calculations reported in Tables 1 and 2. It is clear that while the basic EA-EOMCCSD(2-1) approach can provide qualitative results for the ground states of CH, CNC, and C2N, which are characterized by errors relative to EA-EOMCCSD(3-2) of 2.173, 4.881, and 4.696 millihartree, respectively, it fails to describe the valence excited states, producing errors on the order of 0.1 hartree. The inability of EA-EOMCCSD(2-1) to balance the description of ground and excited states of CH, CNC, and C2N results in highly inaccurate adiabatic excitation spectra, which deviate from their EA-EOMCCSD(3-2) counterparts by 2β4 eV [15, 18]. The failure of EA-EOMCCSD(2-1) is largely a consequence of neglecting the 3-2 effects, which are essential for accurately describing electronic states with significant 2-1 character. This can be mitigated by the noniterative EA-CR-EOMCC(2,3) corrections. As shown in Tables 1 and 2, the 2β5 millihartree differences between the EA-EOMCCSD(2-1) and EA-EOMCCSD(3-2) energies of the ground states of CH, CNC, and C2N are reduced by a factor of 2β3 when the EA-CR-EOMCC(2,3) corrections to EA-EOMCCSD(2-1) are employed. In the case of excited states, the improvements offered by the EA-CR-EOMCC(2,3) corrections are even more substantial. The massive errors relative to EA-EOMCCSD(3-2) obtained with EA-EOMCCSD(2-1), which are 69.816β117.876 millihartree for CH, 112.241β118.827 millihartree for CNC, and 111.733β158.282 millihartree for C2N, reduce in the EA-CR-EOMCC(2,3)A calculations to 17.332β31.850, 20.996β22.060, and 20.003β29.695 millihartree, respectively. The analogous errors characterizing the EA-CR-EOMCC(2,3)D calculations, which are 0.115β12.505 millihartree for CH, 14.346β14.517 millihartree for CNC, and 18.601β33.535 millihartree for C2N, are similar or smaller. The noniterative EA-CR-EOMCC(2,3)A and EA-CR-EOMCC(2,3)D corrections also improve the adiabatic excitation spectra resulting from the EA-EOMCCSD(2-1) calculations, reducing the 2β4 eV errors relative to EA-EOMCCSD(3-2) obtained with EA-EOMCCSD(2-1) to 0.45β0.84 and 0.03β0.97 eV, respectively. While the EA-CR-EOMCC(2,3) methodology helps temper the failures observed in the EA-EOMCCSD(2-1) calculations, the 20β30 millihartree errors relative to EA-EOMCCSD(3-2) characterizing the majority of our EA-CR-EOMCC(2,3) results remain. In analogy to the particle-conserving CC(;) framework [27, 28, 33, 34, 35, 36], this can be attributed to the decoupling of the component of from its lower-rank and counterparts when the EA-CR-EOMCC(2,3) corrections are constructed. As shown in Tables 1 and 2, this problem can be addressed by turning to the EA-CC(t;3) approach, which replaces EA-EOMCCSD(2-1) by EA-EOMCCSDt prior to determining the 3-2 corrections.
@@
Upon examining the EA-EOMCCSDt and EA-CC(t;3) data reported in Tables 1 and 2, the benefits offered by including the dominant 3-2 excitations selected via active orbitals in the iterative EA-EOMCC steps are readily apparent. The energies resulting from the EA-EOMCCSDt calculations for the ground and excited states of CH, CNC, and C2N differ from their EA-EOMCCSD(3-2) parents by 1.454β3.243, 1.414β1.910, and 0.543β1.606 millihartree, respectively, which in most cases is a massive improvement over EA-EOMCCSD(2-1). In fact, with the exception of the and states of CH and the state of CNC, EA-EOMCCSDt outperforms EA-CR-EOMCC(2,3) as well, which highlights the importance of incorporating the coupling between and its lower-rank and counterparts when treating states characterized by stronger 2-1 correlations. The adiabatic excitation spectra of CH, CNC, and C2N computed with EA-EOMCCSDt are also very accurate, with errors relative to full EA-EOMCCSD(3-2) not exceeding 0.05 eV (cf.Β Refs.Β [15, 18]). By correcting the EA-EOMCCSDt energetics for the remaining 3-2 correlations using EA-CC(t;3), we obtain a virtually perfect agreement with the parent EA-EOMCCSD(3-2) data. Indeed, when using the EA-CC(t;3)A approach, the 1β3 millihartree errors relative to EA-EOMCCSD(3-2) resulting from the EA-EOMCCSDt calculations for CH, CNC, and C2N reduce to 0.221β0.412, 0.194β0.241, and 0.084β0.201 millihartree, respectively. The EA-CC(t;3)D corrections to EA-EOMCCSDt, which use the EN denominators in calculating the coefficients in Eq.Β (9) instead of their MP counterparts, are even more accurate, reproducing the target EA-EOMCCSD(3-2) energetics to within tens of microhartree or less. Even the largest deviation of 0.106 millihartree, observed in the EA-CC(t;3)D calculations for the state of CH, is still an order of magnitude smaller than the error obtained with EA-EOMCCSDt. It is important to emphasize that the EA-CC(t;3) calculations reported in Tables 1 and 2 are much less expensive than the parent EA-EOMCCSD(3-2) computations, while providing virtually identical energetics. For example, the spaces adopted in our EA-EOMCCSDt and EA-CC(t;3) computations for CNC and C2N used only about 30% of all 3-2 determinants. In the case of CH, the fractions of 3-2 determinants included in the corresponding spaces were even smaller (less than 5% of all 3-2 determinants). As demonstrated in Ref.Β [15], the use of small fractions of 3-2 determinants in the iterative EA-EOMCCSDt steps translates into substantial savings in the computational costs relative to the parent EA-EOMCCSD(3-2) approach. The EA-CC(t;3) corrections to EA-EOMCCSDt do not change this, since they involve fewer CPU operations than a single iteration of EA-EOMCCSD(3-2).
@@
Similar accuracy patterns are observed in the results of the CC(;)-based IP-EOMCC calculations reported in Tables 3 and 4. Much like EA-EOMCCSD(2-1), the basic IP-EOMCCSD(2-1) approach is generally incapable of providing an accurate description, especially when 2-1 effects become more substantial [15, 18]. In the case of SH, the discrepancies between the IP-EOMCCSD(2-1) and parent IP-EOMCCSD(3-2) energetics are on the order of 0.1β0.2 hartree for all but the and states, which are described more accurately by IP-EOMCCSD(2-1) since these two states are the only ones in Table 3 dominated by 1 correlations. The same is true when examining the vertical excitation spectrum of SH obtained with IP-EOMCCSD(2-1), where the errors relative to IP-EOMCCSD(3-2) are 0.06 eV for the state and 3.5β5.3 eV for the remaining excited states shown in Table 3 (cf.Β Ref.Β [15]). In the case of the electronic states of N3 and NCO reported in Table 4, for which 2-1 effects are small, the IP-EOMCCSD(2-1) method fares better, providing adiabatic transition energies in good agreement (0.3 eV or less) with the IP-EOMCCSD(3-2) results [18], but this is partly due to a fortuitous cancellation of the 10β20 millihartree errors characterizing the IP-EOMCCSD(2-1) energies relative to IP-EOMCCSD(3-2). In analogy to the EA-EOMCC computations, we can improve the IP-EOMCCSD(2-1) results by applying the IP-CR-EOMCC(2,3) corrections due to 3-2 effects. Indeed, as shown in Table 3, the 9.898 and 12.149 millihartree errors relative to IP-EOMCCSD(3-2) obtained with IP-EOMCCSD(2-1) for the and states of SH reduce in the IP-CR-EOMCC(2,3)A calculations to 2.904 and 4.449 millihartree, respectively. The IP-CR-EOMCC(2,3)D approach reduces these errors even further, to 1.576 and 3.201 millihartree. The error reductions offered by the IP-CR-EOMCC(2,3) corrections are even larger when the remaining excited states of SH considered in Table 3, characterized by significant 2-1 contributions, are examined. For these states, the errors relative to IP-EOMCCSD(3-2) obtained with IP-EOMCCSD(2-1), which range from 137.280 to 203.145 millihartree, reduce in the IP-CR-EOMCC(2,3)A and IP-CR-EOMCC(2,3)D calculations to 0.359β39.214 and 3.201β64.470 millihartree, respectively. As a result, the vertical excitation energies corresponding to the state of SH obtained with the IP-CR-EOMCC(2,3) corrections agree with their IP-EOMCCSD(3-2) parent value to within 0.04 eV and the 3.5β5.3 eV differences between the IP-EOMCCSD(2-1) and IP-EOMCCSD(3-2) excitation energies characterizing the remaining excited states of SH shown in Table 3 decrease in the IP-CR-EOMCC(2,3)A and IP-CR-EOMCC(2,3)D computations to 0.07β0.99 eV and 0.14β1.71 eV, respectively. For N3 and NCO, the improvements offered by the 3-2 corrections to IP-EOMCCSD(2-1) are less dramatic, since the electronic states of these two molecules included in Table 4 are dominated by 1 excitations out of (N3)- and (NCO)-, but the reductions in errors relative to IP-EOMCCSD(3-2), from 9.587β20.154 millihartree obtained with IP-EOMCCSD(2-1) to 0.808β7.044 millihartree obtained with variants A and D of IP-CR-EOMCC(2,3), are still quite substantial. We can conclude that the IP-CR-EOMCC(2,3) corrections are helpful, but, as shown in Tables 3 and 4, they are not as effective in approximating the IP-EOMCCSD(3-2) energetics as the IP-CC(t;3) corrections to IP-EOMCCSDt.
@@
The IP-CC(t;3) results reported in Tables 3 and 4 and their IP-EOMCCSDt counterparts clearly show the benefits of including the leading 3-2 determinants in the relevant spaces prior to determining the 3-2 corrections. Indeed, with the exception of the state of NCO, the IP-EOMCCSDt approach substantially improves the IP-EOMCCSD(2-1) energetics, reducing the 9.898β203.145, 13.078β14.623, and 9.587β20.154 millihartree errors relative to IP-EOMCCSD(3-2) obtained for SH, N3, and NCO with IP-EOMCCSD(2-1) to 0.015β1.003, 1.400β6.215, and 1.926β2.205 millihartree, respectively. In the case of the state of NCO, the improvement over the basic IP-EOMCCSD(2-1) level offered by IP-EOMCCSDt is minimal, but this is no longer an issue when the IP-CC(t;3)A and IP-CC(t;3)D corrections to IP-EOMCCSDt are employed. The results in Tables 3 and 4 clearly show that the IP-CC(t;3) corrections not only improve the IP-EOMCCSDt energetics, but they are also considerably more accurate than their IP-EOMCCSD(2-1)-based IP-CR-EOMCC(2,3) counterparts, and this applies to all of the electronic states of SH, N3, and NCO considered in this work. For SH, both IP-CC(t;3)A and IP-CC(t;3)D recover the parent IP-EOMCCSD(3-2) data to within small fractions of a millihartree. In the cases of N3 and NCO, the IP-CC(t;3)A method shows a maximum error relative to IP-EOMCCSD(3-2) of just 1.690 millihartree, while the more complete IP-CC(t;3)D approach recovers its parent energetics to within one millihartree for all electronic states. In analogy to the EA-CC(t;3) approach, the IP-CC(t;3) calculations provide IP-EOMCCSD(3-2)-level results with reduced computational costs. Indeed, the spaces entering our IP-EOMCCSDt and IP-CC(t;3) calculations for SH, N3, and NCO include approximately 60% of all 3-2 determinants, which translate to sizeable speedups compared to the full IP-EOMCCSD(3-2) computations. In fact, the iterative IP-EOMCCSDt and noniterative IP-CC(t;3) steps require CPU efforts that are comparable to the underlying -electron CCSD calculations.

### v2

@@ Abstract
1 Introduction
@@ 2 Theory
This study focuses on the CC(;) approaches aimed at accurately approximating the EA-EOMCCSD(3-2) and IP-EOMCCSD(3-2) energetics at small fractions of the computational costs by correcting the results of the EA/IP-EOMCC calculations with all 1/1, all 2-1/2-1, and, optionally, some 3-2/3-2 excitations on top of CCSD for the many-electron correlation effects associated with all or the remaining 3-2/3-2 contributions using the moment expansions defined by Eq.Β (9). In the case of the noniterative corrections to EA-EOMCCSD(2-1) due to 3-2 correlations, which define the EA-CR-EOMCC(2,3) method, the space is spanned by all 1 () and 2-1 () determinants and the corresponding space consists of all 3-2 determinants (). Similarly, the IP-CR-EOMCC(2,3) approach, which corrects the IP-EOMCCSD(2-1) energetics for 3-2 effects, uses a space spanned by all 1 () and 2-1 () determinants and a space consisting of all 3-2 determinants (). In the case of the active-orbital-based EA-CC(t;3) methodology, the space is defined as all and determinants and the subset of 3-2 determinants selected according to the formula , where the index runs over active unoccupied spinorbitals defined by the user, and the space consists of the remaining 3-2 determinants that are not included in the space. Finally, the IP-CC(t;3) approach adopts a space spanned by all , all , and the subset of 3-2 determinants defined as , where the index denotes the user-defined active occupied spinorbitals, and the space is spanned by the remaining 3-2 determinants. It is worth mentioning that one could extend the EA/IP-based CC(;) methods to capture the correlation effects due to all or a subset of 4-3/4-3, 5-4/5-4, etc.Β excitations, in addition to all or some 3-2/3-2 determinants, provided that the truncation of the -electron cluster operator, , is greater than or equal to the highest value of characterizing the -/- determinants included in and (a condition required to retain size-intensivity [13, 14, 15]). For example, we could perform CC(;) calculations using an ()-electron space spanned by all 1/1, 2-1/2-1, and a subset of 3-2/3-2 and 4-3/4-3 excitations, and correct the results for the missing 3-2/3-2 and 4-3/4-3 effects, if we used CCSDT () to describe the -electron reference species.
@@
In analogy to the CC(;) methods of Refs.Β [27, 28, 33, 34, 35, 36] and their CR-CC(2,3) and CR-EOMCC(2,3) predecessors, the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches offer considerable savings in the computational effort compared to their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents. This is because the excitation spaces used in the iterative EA/IP-EOMCC steps contain none or only some 3-2/3-2 determinants and the noniterative corrections require fewer CPU operations than a single iteration of EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) (cf.Β Table 1 for the computational costs characterizing the different EA/IP-EOMCC methods considered in this work). In the next section, we assess how effective the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) methods are in improving the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt computations and recovering the parent EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) energetics. In doing so, we examine two variants of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, termed EA/IP-CR-EOMCC(2,3)X and EA/IP-CC(t;3)X, respectively, where or . Following the naming convention adopted in Refs.Β [31, 32, 33], the methods labeled with subscript use the EN-like energy denominators in Eq.Β (9) when evaluating corrections . The methods using subscript replace the EN form of by its MP counterpart.
@@ 3.1 Computational Details
To assess the performance of the EA/IP-CR-EOMCC(2,3) and EA/IP-CC(t;3) approaches, we applied them, along with the underlying EA-EOMCCSD(2-1)/IP-EOMCCSD(2-1) and EA/IP-EOMCCSDt methods and their EA-EOMCCSD(3-2)/IP-EOMCCSD(3-2) parents, to the ground and valence excited states of the CH, CNC, C2N, SH, N3, and NCO radical species. The results of our EA-EOMCC calculations for CH, CNC, and C2N are reported in Tables 2 and 3, whereas Tables 4 and 5 summarize the IP-EOMCC computations for SH, N3, and NCO. With the exception of SH, all of our EA/IP-EOMCC calculations correspond to adiabatic transitions; calculations for SH were performed at the ground-state equilibrium geometry. The nuclear geometries characterizing the ground and excited states of CH and the ground-state SβH separation in SH were the same as those used in Ref.Β [15], whereas the CNC, C2N, N3, and NCO molecules were described using the linear geometries optimized with the EA (CNC and C2N) or IP (N3 and NCO) variants of the symmetry-adapted-cluster (SAC) configuration interaction (CI) approach [38] abbreviated as SDT- [39] in Ref.Β [18]. All EA-EOMCC calculations for CH, CNC, and C2N reported in this work employed the restricted HartreeβFock (RHF) orbitals of the underlying (CH)+, (CNC)+, and (C2N)+ cationic references, whereas the IP-EOMCC calculations for SH, N3, and NCO used the RHF orbitals of the closed-shell (SH)-, (N3)-, (NCO)- anions. Our calculations for CH employed the aug-cc-pVQZ basis set [40, 41], while for SH, we used the aug-cc-pV(T+d)Z [42] basis for sulfur and aug-cc-pVTZ [40, 41] for hydrogen. Following Ref.Β [18], the CNC, C2N, N3, and NCO radicals were described using the DZP[4s2p1d] basis set [43, 44]. The active orbital spaces used to define the and components of entering the EA/IP-EOMCCSDt and subsequent EA/IP-CC(t;3) calculations were the same as those used in Refs.Β [15, 18]. Thus, the EA-EOMCCSDt and EA-CC(t;3) calculations for CH used an active space consisting of the lowest-energy orbitals unoccupied in (CH)+, whereas the active spaces employed in the analogous calculations for CNC and C2N were spanned by the four orbitals corresponding to the two lowest shells unoccupied in (CNC)+ and (C2N)+. In the case of the IP-EOMCCSDt and IP-CC(t;3) computations for SH, N3 and NCO, the active spaces consisted of the highest-energy shell occupied in the respective parent (SH)-, (N3)-, and (NCO)- anions. The lowest-energy orbitals corresponding to the 1s shells of carbon, oxygen, nitrogen, and sulfur atoms were frozen in all post-RHF steps.
@@ 3.2 Performance of the CC(;)-based EA/IP-EOMCC Methodologies
We begin our discussion by examining the EA-EOMCC calculations reported in Tables 2 and 3. As previously shown in Refs.Β [15, 18], the basic EA-EOMCCSD(2-1) method can only treat the ground states of CH, CNC, and C2N with reasonable accuracy relative to EA-EOMCCSD(3-2) (to within 5 millihartree), but it breaks down when describing the valence excited states, producing errors on the order of 0.1 hartree, which lead to adiabatic excitation energies that deviate from EA-EOMCCSD(3-2) by 2β4 eV. The failure of EA-EOMCCSD(2-1) is largely a consequence of neglecting the 3-2 effects, which are essential for accurately describing electronic states with significant 2-1 character. This can be mitigated by the noniterative EA-CR-EOMCC(2,3) corrections. As shown in Tables 2 and 3, the 2β5 millihartree differences between the EA-EOMCCSD(2-1) and EA-EOMCCSD(3-2) energies of the ground states of CH, CNC, and C2N are reduced by a factor of 2β3 when the EA-CR-EOMCC(2,3) corrections to EA-EOMCCSD(2-1) are employed. In the case of excited states, the improvements offered by the EA-CR-EOMCC(2,3) corrections are even more substantial. The massive errors of EA-EOMCCSD(2-1) relative to EA-EOMCCSD(3-2), which usually exceed 100 millihartree, are significantly reduced to 17.332β31.850 millihartree for CH, 20.996β22.060 millihartree for CNC, and 20.003β29.695 millihartree for C2N when using EA-CR-EOMCC(2,3)A. The analogous errors characterizing the EA-CR-EOMCC(2,3)D calculations, which are 0.115β12.505 millihartree for CH, 14.346β14.517 millihartree for CNC, and 18.601β33.535 millihartree for C2N, are similar or smaller. The noniterative EA-CR-EOMCC(2,3)A and EA-CR-EOMCC(2,3)D corrections also improve the adiabatic excitation spectra resulting from the EA-EOMCCSD(2-1) calculations, reducing the 2β4 eV errors relative to EA-EOMCCSD(3-2) obtained with EA-EOMCCSD(2-1) to 0.45β0.84 and 0.03β0.97 eV, respectively. While the EA-CR-EOMCC(2,3) methodology helps temper the failures of EA-EOMCCSD(2-1), the 20β30 millihartree errors relative to EA-EOMCCSD(3-2) characterizing the majority of our EA-CR-EOMCC(2,3) results remain. In analogy to the particle-conserving CC(;) framework [27, 28, 33, 34, 35, 36], this can be attributed to the decoupling of the component of from its lower-rank and counterparts when the EA-CR-EOMCC(2,3) corrections are constructed. As shown in Tables 2 and 3, this problem can be addressed by turning to the EA-CC(t;3) approach, which replaces EA-EOMCCSD(2-1) by EA-EOMCCSDt prior to determining the 3-2 corrections.
@@
Upon examining the EA-EOMCCSDt and EA-CC(t;3) data reported in Tables 2 and 3, the benefits offered by including the dominant 3-2 excitations selected via active orbitals in the iterative EA-EOMCC steps are readily apparent. The energies resulting from the EA-EOMCCSDt calculations for the ground and excited states of CH, CNC, and C2N differ from their EA-EOMCCSD(3-2) parents by 1.454β3.243, 1.414β1.910, and 0.543β1.606 millihartree, respectively, which in most cases is a massive improvement over EA-EOMCCSD(2-1). In fact, with the exception of the and states of CH and the state of CNC, EA-EOMCCSDt outperforms EA-CR-EOMCC(2,3) as well, which highlights the importance of incorporating the coupling between and its lower-rank and counterparts when treating states characterized by stronger 2-1 correlations. The adiabatic excitation spectra of CH, CNC, and C2N computed with EA-EOMCCSDt are also very accurate, with errors relative to full EA-EOMCCSD(3-2) not exceeding 0.05 eV (cf.Β Refs.Β [15, 18]). By correcting the EA-EOMCCSDt energetics for the remaining 3-2 correlations using EA-CC(t;3), we obtain a virtually perfect agreement with the parent EA-EOMCCSD(3-2) data. Indeed, when using the EA-CC(t;3)A approach, the 1β3 millihartree errors relative to EA-EOMCCSD(3-2) resulting from the EA-EOMCCSDt calculations for CH, CNC, and C2N reduce to 0.221β0.412, 0.194β0.241, and 0.084β0.201 millihartree, respectively. The EA-CC(t;3)D corrections to EA-EOMCCSDt are even more accurate, reproducing the target EA-EOMCCSD(3-2) energetics to within tens of microhartree or less. Even the largest deviation of 0.106 millihartree, observed in the EA-CC(t;3)D calculations for the state of CH, is still an order of magnitude smaller than the error obtained with EA-EOMCCSDt. It is important to emphasize that the EA-CC(t;3) calculations reported in Tables 2 and 3 are much less expensive than the parent EA-EOMCCSD(3-2) computations, while providing virtually identical energetics. For example, the spaces adopted in our EA-EOMCCSDt and EA-CC(t;3) computations for CNC and C2N used only about 30% of all 3-2 determinants. In the case of CH, the fractions of 3-2 determinants included in the corresponding spaces were even smaller (less than 5% of all 3-2 determinants). As demonstrated in Ref.Β [15], the use of small fractions of 3-2 determinants in the iterative EA-EOMCCSDt steps translates into substantial savings in the computational costs relative to the parent EA-EOMCCSD(3-2) approach. The EA-CC(t;3) corrections to EA-EOMCCSDt do not change this, since they involve fewer CPU operations than a single iteration of EA-EOMCCSD(3-2).
@@
Similar accuracy patterns are observed in the results of the CC(;)-based IP-EOMCC calculations reported in Tables 4 and 5. Much like EA-EOMCCSD(2-1), the basic IP-EOMCCSD(2-1) approach is generally incapable of providing an accurate description, especially when 2-1 effects become more substantial [15, 18]. In the case of SH, the discrepancies between the IP-EOMCCSD(2-1) and parent IP-EOMCCSD(3-2) energetics are on the order of 0.1β0.2 hartree for all but the and states, which are described more accurately by IP-EOMCCSD(2-1) since these two states are the only ones in Table 4 dominated by 1 correlations. The same is true when examining the vertical excitation spectrum of SH obtained with IP-EOMCCSD(2-1), where the errors relative to IP-EOMCCSD(3-2) are 0.06 eV for the state and 3.5β5.3 eV for the remaining excited states shown in Table 4 (cf.Β Ref.Β [15]). In the case of the electronic states of N3 and NCO reported in Table 5, for which 2-1 effects are small, the IP-EOMCCSD(2-1) method fares better, providing adiabatic transition energies in good agreement (0.3 eV or less) with the IP-EOMCCSD(3-2) results [18], but this is partly due to a fortuitous cancellation of the 10β20 millihartree errors characterizing the IP-EOMCCSD(2-1) energies relative to IP-EOMCCSD(3-2). In analogy to the EA-EOMCC computations, we can improve the IP-EOMCCSD(2-1) results by applying the IP-CR-EOMCC(2,3) corrections due to 3-2 effects. Indeed, as shown in Table 4, the errors characterizing the IP-EOMCCSD(2-1) results for the and states of SH decrease by a factor of nearly three or more using either variant of the IP-CR-EOMCC(2,3) approach. The error reductions offered by the IP-CR-EOMCC(2,3) corrections are even larger when examining the remaining excited states of SH considered in Table 4, which are characterized by significant 2-1 contributions. For these states, the 137.280β203.145 millihartree errors relative to IP-EOMCCSD(3-2) obtained with IP-EOMCCSD(2-1) reduce in the IP-CR-EOMCC(2,3)A and IP-CR-EOMCC(2,3)D calculations to 0.359β39.214 and 3.201β64.470 millihartree, respectively. As a result, the vertical excitation energies of SH obtained with IP-CR-EOMCC(2,3) greatly improve upon IP-EOMCCSD(2-1), agreeing with their IP-EOMCCSD(3-2) parent values to within 0.04β1.71 eV. For N3 and NCO, the improvements offered by the 3-2 corrections to IP-EOMCCSD(2-1) are less dramatic, since the electronic states of these two molecules included in Table 5 are dominated by 1 excitations out of (N3)- and (NCO)-, but the reductions in errors relative to IP-EOMCCSD(3-2), from 9.587β20.154 millihartree obtained with IP-EOMCCSD(2-1) to 0.808β7.044 millihartree obtained with variants A and D of IP-CR-EOMCC(2,3), are still quite substantial. We can conclude that the IP-CR-EOMCC(2,3) corrections are helpful, but, as shown in Tables 4 and 5, they are not as effective in approximating the IP-EOMCCSD(3-2) energetics as the IP-CC(t;3) corrections to IP-EOMCCSDt.
@@
The IP-CC(t;3) results reported in Tables 4 and 5 and their IP-EOMCCSDt counterparts clearly show the benefits of including the leading 3-2 determinants in the relevant spaces prior to determining the 3-2 corrections. Indeed, with the exception of the state of NCO, the IP-EOMCCSDt approach substantially improves the IP-EOMCCSD(2-1) energetics, reducing the 9.898β203.145, 13.078β14.623, and 9.587β20.154 millihartree errors relative to IP-EOMCCSD(3-2) obtained for SH, N3, and NCO with IP-EOMCCSD(2-1) to 0.015β1.003, 1.400β6.215, and 1.926β2.205 millihartree, respectively. In the case of the state of NCO, the improvement over the basic IP-EOMCCSD(2-1) level offered by IP-EOMCCSDt is minimal, but this is no longer an issue when the IP-CC(t;3)A and IP-CC(t;3)D corrections to IP-EOMCCSDt are employed. The results in Tables 4 and 5 clearly show that the IP-CC(t;3) corrections not only improve the IP-EOMCCSDt energetics, but they are also considerably more accurate than their IP-EOMCCSD(2-1)-based IP-CR-EOMCC(2,3) counterparts, and this applies to all of the electronic states of SH, N3, and NCO considered in this work. For SH, both IP-CC(t;3)A and IP-CC(t;3)D recover the parent IP-EOMCCSD(3-2) data to within small fractions of a millihartree. In the cases of N3 and NCO, the IP-CC(t;3)A method shows a maximum error relative to IP-EOMCCSD(3-2) of just 1.690 millihartree, while the more complete IP-CC(t;3)D approach recovers its parent energetics to within one millihartree for all electronic states. In analogy to the EA-CC(t;3) approach, the IP-CC(t;3) calculations provide IP-EOMCCSD(3-2)-level results with reduced computational costs. Indeed, the spaces entering our IP-EOMCCSDt and IP-CC(t;3) calculations for SH, N3, and NCO include approximately 60% of all 3-2 determinants, which translate to sizeable speedups compared to the full IP-EOMCCSD(3-2) computations. In fact, the iterative IP-EOMCCSDt and noniterative IP-CC(t;3) steps require CPU efforts that are comparable to the underlying -electron CCSD calculations.